A voltage V across a resistance R generates a current I =
V/R. If a constant voltage of 22 volts is put across a resistance
that is increasing at a rate of 0.2 ohms per second when the re￾sistance is 5 ohms, at what rate is the current changing?
1
Expert's answer
2020-03-09T14:13:59-0400

The relation among voltage "(V)" , resistance "(R)" and current "(I)" is given by the following equation:

"I = \frac{V}{R} \hspace{1 cm}" (1)

We are given:

The constant votage is:

"V = 22" volts

The resistance is:

"R = 5" ohms.

And:

The increasing rate of resistance is:

"\frac{dR}{dt} = 0.2" ohms per second

Now:

The derivative of equation(1) with respect to time "t" is:

"\frac{dI}{dt} = V \frac{d}{dt} (\frac{1}{R})" "\hspace{1 cm}" [ Because V is a constant ]

"\,\,\,\,\,\,\,\,= V \frac{d}{dt} (R^{-1} )"

"\,\,\,\,\,\,\,\,= V \left( - \frac{1}{R^2} \frac{dR}{dt} \right)" "\hspace{1 cm} \left[ \because \frac{d}{dx} (x^n) = nx^{n-1} \right]"

"\,\,\,\,\,\,\,\,= - \frac{22}{5^2} .(0.2)" "\hspace{1cm} \left[ \because V = 22 \, , \, R = 5 \, and \, \frac{dR}{dt} = 0.2 \right]"

"\,\,\,\,\,\,\, \, = - 0.176"

Hence:

The decreasing rate of current is "0.176" amperes per second